OpenStudy (anonymous):

3x+12/4x-18 divided by x+4/x-3

OpenStudy (anonymous):

dividing by a fraction is the same as multiplying by its reciprocal ex. \[\frac{a}{b} \div \frac{c}{d} = \frac{ a}{b} \times \frac{d}{c}\]

OpenStudy (anonymous):

3(x+4) x-3 3(x-3) ------- * ------ = ------- 2(2x-9) x+4 2(2x-9)

OpenStudy (anonymous):

@Kimmieboo1316